# CrowdGenUI: Enhancing LLM-Based UI Widget Generation with a Crowdsourced Preference Library

### v1

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable skills across various design domains, including UI generation. However, current LLMs for UI generation tend to offer generic solutions that lack a deep understanding of task context and user preferences in specific scenarios. We present CrowdGenUI, a framework that enhances LLM-driven UI generation with a crowdsourced user preference library. This approach addresses the limitations of existing methods by guiding LLM reasoning with user preferences, enabling the generation of UI widgets that align more closely with user needs and task-specific requirements. Using image editing as a test domain, we built this library from 50 users, capturing 720 user preferences, which include the predictability, efficiency, and explorability of multiple UI widgets. In a user study with 72 additional participants, our framework outperformed standard LLM-generated widgets in meeting user preferences and task requirements. We discuss these findings to inform future opportunities for designing user-centered and customizable UIs by comprehensively analyzing the extendability of the proposed framework and crowdsourced library.
@@
$c$s: Human-centered computing Collaborative and social computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing User studiesccs: Information systems Crowdsourcing
Figure 1. CrowdGenUI is a framework that allows the generation of task requirement-aware and user preference-aligned UI widgets powered by LLMs. Based on input information specifying the content to edit, the task to perform, and various user preferences, the LLM-driven UI widget generation is enhanced by our crowdsourced user preference library together with a set of widget candidates for UI widget reasoning and code generation.
@@ 1. Introduction
Graphical user interfaces (GUIs) play a fundamental role in how most people operate computers and other computing devices. Carefully designed by expert interface designers, these GUIs aim to provide users with learnable ways to interact with software (Adobe, 2024c, b). Although effective, such interfaces can be complicated and sometimes overwhelming for users. In recent years, the emergence of large language models (LLMs) has led to increased interest in natural language interfaces. These language-based interfaces process human language input, analyze high-level intentions, and turn the interpreted intents into operations, providing a simplified way for users to interact with technology (OpenAI, 2024a; Descript, 2024; Maddigan and Susnjak, 2023). However, while language interfaces are promising, they are typically suited for high-level task execution, such as removing an object from an image or adding a filtering effect to videos. They often lack the precision required when users need to fine-tune specific parameters, such as adjusting the color in an image or customizing detailed settings in a video editing project.
To address the complexity of professional GUIs and the inaccuracy of natural language interfaces, researchers have begun to explore the integration of LLMs with UI generation to combine the high-level task understanding enabled by LLMs with the precise, granular control offered by GUIs (Vaithilingam et al., 2024; Cheng et al., 2024). By turning natural language inputs into interactable UI widgets through UI generation, these systems provide an exciting hybrid interaction model that supports both user interaction with widgets for precise control and language as input for high-level task interpretation. Nevertheless, the existing UI generation systems also come with challenges. These LLM-generated UIs tend to be generic and may not fully account for the nuances of domain-specific tasks or the diverse preferences of individual users. Without specialized knowledge of interface design or a deeper understanding of user preferences, these UIs may fail to provide the optimal interaction experience.
@@
To overcome these challenges, we propose a novel framework, CrowdGenUI, that enhances LLM-driven UI generation by incorporating a crowdsourced user preference library, as illustrated in Figure 1. This framework leverages the collective insights of users to guide the LLM in generating tailored UI widgets for different user preferences. While the framework is broadly applicable across various domains, we demonstrate its effectiveness in the context of image editing within a Python programming environment. To build the preference library, we collected data from 50 users, with no requirement on UI design or programming experience, capturing their preferences across three key dimensions of interface design, i.e., predictability, efficiency, and explorability. By integrating this library into the LLM’s UI widget generation process, our system generates widgets that are not only more precise but also better aligned with user preferences and task-specific requirements.
@@
In our user study with a different set of 72 participants, we found that UI widgets generated based on our proposed framework were more aligned with user preferences and task requirements than those generic widgets generated by the LLM alone. This demonstrates the potential of our framework to improve LLM-based UI generation and suggests that crowdsourced user preference data can play a crucial role in creating user-centered user interfaces.
To summarize, this paper makes the following key contributions:
  • We propose a general framework to enhance LLM-driven UI generation by incorporating a crowdsourced user preference library for task-aware and preference-aligned widget generation.
  • As a concrete implementation of this framework, we develop a system for the domain of image editing within a Python programming environment. This system includes crowdsourced data from 50 users, capturing user preferences for UI widget configurations in image editing tasks.
  • We conduct a user study with an additional 72 participants to evaluate the effectiveness of the implemented system. The results of this study are extensively discussed to inform future research on the development of customizable and user-centered UIs.
@@ 2. Related Work
To provide the background for our work, we introduce prior work on traditional fixed UI, natural language UI, and dynamically generated UI. We discuss their advantages and limitations and compare them with our proposed framework. A visual comparison is presented in Figure 2.
2.1. Traditional Fixed UI
@@
Traditional UI design in professional digital software like Illustrator (Adobe, 2024b), Photoshop (Adobe, 2024c), Premiere Pro (Adobe, 2024e), and Blender (Blender, 2024) follows a designer-centric model, where a fixed set of tools and controls are designed for various tasks. These applications rely on well-established interface elements such as dropdown menus, sliders, text fields, and panels that allow users to manipulate images, graphics, videos, and 3D models. For example, Photoshop provides consistent tools for adjustments like brightness, contrast, and hue, while Illustrator offers vector editing through pre-defined tools. Premiere Pro delivers video editing features through timelines and control panels, and Blender offers a complex interface for 3D modeling and animation. These tools are powerful but often rigid, requiring users to adapt to the software’s structure. As a result, their complexity can be overwhelming to users, and their interface may not always align with individual workflows or preferences, which limits the flexibility and personalization of user experience.
$2.$ Natural Language UI
Using natural language as input to edit digital content has gained increasing interest due to its simplicity. This approach allows users to perform tasks like image, video, data visualization, 3D models, and human motion editing by simply describing the changes they want rather than navigating complex software interfaces. For example, natural language-based image editing tools, such as DALL-E (OpenAI, 2024a) or Runway (Runaway, 2024), allow users to generate and edit images by typing commands like “make the sky bluer” or “add a tree in the background”. Similarly, tools like Adobe Firefly (Adobe, 2024a) and Photoshop Generative Fill (Adobe, 2024d) enable image generation and edits using prompts like “change to a vintage look” or “add a yellow bee in the circled region next to the flower”. In video editing, tools like Descript (Descript, 2024) allow users to edit video content by editing a transcript, and tools like Runway and Premiere Pro with Sensei AI (Adobe, 2024f) also support video editing using natural language prompts like “add a motion blur effect”. Also, there are many research projects (Brooks et al., 2023; Zhang et al., 2024; Fu et al., 2024, 2022; Qin et al., 2024) that have significantly advanced the techniques for image and video editing using language as the primary modality. In the field of visualization, systems like Chat2VIS (Maddigan and Susnjak, 2023), NL4DV (Sah et al., 2024) and Data Formulator (Wang et al., 2023, 2024) allow users to create charts and graphs by simply asking for a “bar chart showing sales over time”, and LIDA (Dibia, 2023) enable the generation of grammar-agnostic visualizations and infographics using LLMs. For the generation and editing of 3D model (Chang et al., 2017; Poole et al., 2022) and human motion (Tevet et al., 2022; Raab et al., 2023; Chen et al., 2023; Liu and Sra, 2024), prior research has investigated using language to describe users’ goals, such as “create a 3D lemur taking notes in a journal” or “change the upper body motion to throwing a ball and keep the original running motion”. These innovations are transforming traditional workflows by making advanced editing more accessible to non-expert users, reducing the technical execution gulf and enabling them to focus on creative intent.
However, natural language interfaces struggle with precise control over specific details and cannot effectively support interactive editing, limiting their ability to handle complex, fine-grained adjustments typically required in professional workflows.
2.3. Dynamically Generated UI
@@
Recent research has studied generating dynamic UIs to provide users with precise, direct interactions for content editing. One prominent approach is the relaxation method, which involves creating UI widgets that provide generalized control over variables within a function or query. For instance, Mavo (Verou et al., 2016) provides users with the capability to construct interactive HTML pages by incorporating unique attributes, which, in turn, recommend editing widgets tailored to those attributes. Heer et al. (Heer et al., 2008) employ query relaxation techniques, allowing users to broaden their selection effectively. Bespoke (Vaithilingam and Guo, 2019) utilizes user demonstrations to generate custom UIs for command-line applications. It applies rule-based heuristics to infer the semantic types of parameters in bash commands and thus dynamically create widgets for parameter manipulation. Another line of research, known as precision interfaces, uses SQL queries as a framework for UI generation (Zhang et al., 2018; Chen, 2020; Chen and Wu, 2022). These systems convert a sequence of input queries into interactive widgets, allowing users to adjust query parameters directly. An example is NL2Interface (Chen et al., 2022), which translates natural language commands into SQL queries and subsequently generates a UI for users to edit parameters within the queries.
@@
In addition to generating dynamic UIs from command-line applications and queries, recent advancements have focused on creating UIs directly from natural language descriptions. This approach allows users to articulate their goals in natural language, and the system generates an interactive UI tailored to the task at hand using LLM’s reasoning capabilities. For instance, DynaVis (Vaithilingam et al., 2024) generates dynamic interfaces for visualization editing based on natural language commands. By enabling users to describe changes to their visualizations through simple text, such as “rotate the x-labels by 30 degrees counter-clockwise”, DynaVis dynamically creates suitable widgets, such as a slider for adjusting the rotation of x-labels. Similarly, Biscuit (Cheng et al., 2024) generates UIs to help users follow machine learning tutorials in Jupyter Notebooks using natural language queries. Users can describe tasks like “sample training data” or “adjust model hyperparameters”, and Biscuit generates ephemeral UIs for the specified parameter adjustments.
Although natural language-based UI generation generates UIs without depending on fixed rules or query relaxation techniques and utilizes LLMs’ advantage to interpret intents from high-level descriptions, most prior work using standalone LLMs does not fully consider individual user preferences and domain-specific task requirements. This can result in interfaces that, while functional, may not be the most effective or usable for specific users, especially when personalization and task-specific requirements are needed.
3. User Feedback Collection and Principles for UI Design
While LLM-powered UI generation offers many advantages, it primarily relies on generic training data, lacking task-specific contexts and user-centered insights, such as individual preferences. User preference is part of user feedback and plays a crucial role in UI design and UI widget selection. The user feedback data is traditionally collected through labor-intensive methods like interviews, usability tests, A/B tests, and feedback forms (Stone et al., 2005; Wilson, 2013). Given that LLMs are trained on large-scale human data, they provide a promising solution to incorporating user preferences into dynamically generated UIs. However, without dedicated user preference datasets, current LLMs cannot fully address this need. To fill this gap, we propose crowdsourcing a preference library to enhance LLM-driven UI widget generation with user preference data, motivated by the effectiveness of adopting crowdsourcing for user feedback collection to assist various design areas (Xu et al., 2014, 2015; Luther et al., 2015; Yu et al., 2016; Foong et al., 2017; Lee et al., 2018; Oppenlaender et al., 2020).
To determine the specific aspects of user preference, we refer to well-established UI design principles developed by prior research. McKay’s UI is Communication (McKay, 2013) emphasizes that an intuitive interface relies on effective communication that balances multiple principles, including predictability, efficiency, discoverability, and explorability. McKay argues that intuitive UIs align with users’ prior experiences, providing clear affordances and responsive feedback to guide interaction. Lidwell et al.’s Universal Principles of Design (Lidwell et al., 2010) expand on these ideas by highlighting fundamental design principles such as the flexibility-usability tradeoff, mapping, and mental models, which underscore the importance of aligning system controls with user expectations and simplifying complex interactions. Together, these principles highlight that successful UIs are predictable and efficient and allow for exploratory interaction without overwhelming the user. Drawing from these foundational principles, we have identified predictability, efficiency, and explorability as three critical aspects to guide our approach to building a user preference library to enhance LLM-powered UI widget generation.
The remainder of the paper is organized as follows: first, we provide an overview of our proposed framework. Next, we describe the implemented system in the image editing domain and discuss the crowdsourced data collection process. This is followed by a detailed user evaluation, along with a discussion of the limitations, potential future work, and the conclusion.
@@ 3. User Feedback Collection and Principles for UI Design
Figure 2. Overview of our proposed CrowdGenUI framework and comparison with prior work. Our framework consists of four main components: it takes (1) task description in natural language as input, and this information is passed to the LLM together with (2) the collected user preference library and (3) a collection of UI widget candidates. All the input data is used to guide the LLM reasoning process to obtain (4) generated UI widgets that are aligned with user preferences and task requirements.
4. Framework Overview
Figure 2 provides an overview of our proposed framework, CrowdGenUI, and a visual comparison with previous research. Our framework improves existing UI widget generation systems that take task descriptions as input for LLMs to generate UIs dynamically. To align generated UI widgets with specific user needs, we collect a user preference library that captures users’ preferred UI widgets and their rationale. This library is provided to the LLM alongside task descriptions and widget candidates to guide the UI generation process. For implementation, the widget candidates are selected from the widget list of Jupyter Widgets Library (Jupyter, 2024b). These candidates can be easily extended to include additional widget options, such as those from other widget libraries or LLM-generated widgets.
5. Crowdsourced Library-enhanced UI Widget Generation for Image Editing
Figure 3 illustrates the system we implemented for our proposed framework. In this section, we dive into the technical details of this system, including UI widget reasoning, integration of the crowdsourced library into the reasoning process, and UI widget code generation.

### v2

# CrowdGenUI: Aligning LLM-Based UI Generation with Crowdsourced User Preferences

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable potential across various design domains, including user interface (UI) generation. However, current LLMs for UI generation tend to offer generic solutions that lack a nuanced understanding of task context and user preferences. We present CrowdGenUI, a framework that enhances LLM-based UI generation with crowdsourced user preferences. This framework addresses the limitations by guiding LLM reasoning with real user preferences, enabling the generation of UI widgets that reflect user needs and task-specific requirements. We evaluate our framework in the image editing domain by collecting a library of 720 user preferences from 50 participants, covering preferences such as predictability, efficiency, and explorability of various UI widgets. A user study (N=78) demonstrates that UIs generated with our preference-guided framework can better match user intentions compared to those generated by LLMs alone, highlighting the effectiveness of our proposed framework. We further discuss the study findings and present insights for future research on LLM-based user-centered UI generation.
@@
ccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Graphical user interfacesccs: Human-centered computing User studiesccs: Information systems Crowdsourcing
@@ 1. Introduction
Graphical user interfaces (GUIs) have long served as the primary medium through which users interact with software applications and digital systems (Adobe, 2024c, b). Despite their ubiquity and effectiveness, designing intuitive and efficient GUIs remains a complex and labor-intensive process. It often requires expert knowledge, iterative user testing, and significant time investments—both from designers to refine the UI and from users to learn how to use it effectively.
In response, there has been growing interest in natural language interfaces, which offer an alternative mode of interaction by allowing users to express their intentions through language. These interfaces reduce the design and learning overhead, simplifying how users interact with complex tools (OpenAI, 2024a; Descript, 2024; Maddigan and Susnjak, 2023). However, while convenient, natural language interfaces often fall short in scenarios requiring precision and control, such as adjusting image brightness or fine-tuning the parameters of a filter. Moreover, they typically lack the interactive, real-time feedback that GUIs provide, which is essential for tasks that benefit from trial-and-error and visual confirmation.
With recent advances in generative AI, particularly the emergence of large language models (LLMs), researchers have started exploring how LLMs might bridge this gap by automatically generating adaptive user interfaces (Vaithilingam et al., 2024; Cheng et al., 2024). These systems aim to combine the high-level task understanding and domain flexibility of LLMs with the precise, user-driven control of traditional GUIs. Yet, current LLM-based UI generation methods rely mainly on the model’s internal reasoning and general knowledge. As a result, the generated UIs often appear generic and may fail to capture the nuanced requirements of specific tasks or the varied preferences of individual users.
@@
To address these limitations, we present CrowdGenUI, a novel framework that enhances LLM-based UI generation by incorporating crowdsourced user preferences. This framework leverages collective insights from real users to guide the LLM in generating UI widgets that are better tailored to different user preferences and task contexts. Although broadly applicable, we focus on the image editing domain within a Python-based environment to demonstrate the framework’s effectiveness. We constructed a preference library by collecting data from 50 participants to capture general users’ preferences across three core dimensions of UI design: predictability, efficiency, and explorability. This preference data is then integrated into the LLM’s UI generation process to augment LLM-based widget reasoning and code generation to produce UI widgets that align with user and task demands.
@@
In a user study with 78 additional participants, we found that UI widgets generated using our framework were rated as more aligned with user preferences and task requirements compared to those generated by the LLM alone. These results highlight the potential of our framework to advance user-centered interface generation and demonstrate that integrating crowdsourced preferences can significantly enhance the effectiveness of LLM-based UI generation.
In summary, our main contributions are as follows:
  • We present a framework that enhances LLM-based UI generation by incorporating crowdsourced user preferences to support task-aware and preference-aligned UI widget generation. We introduce the framework design and the methodology for crowdsourcing user preference data.
  • We implement a system prototype for this framework in the image editing domain. This implementation delivers 720 user preferences crowdsourced from 50 users and a structured reasoning technique to guide LLM reasoning and code generation for user preference-aligned UI generation.
  • We conduct a user study with an additional 78 participants that discovers the impact of crowdsourced user preference data on UI generation. By discussing the study findings, we provide insights for future research on creating adaptive and user-centered LLM-generated UIs with user data as guidance.
@@ 2. Related Work
To provide background for our work, we introduce prior research on traditional fixed UIs, natural language UIs, and dynamically generated UIs, and discuss their advantages as well as limitations.
2.1. Traditional Fixed UIs
@@
Traditional UI design in professional digital software like Illustrator (Adobe, 2024b), Photoshop (Adobe, 2024c), Premiere Pro (Adobe, 2024e), and Blender (Blender, 2024) follows a designer-centric model, where a fixed set of tools and UI controls are designed for various tasks. These applications rely on well-established interface elements such as dropdown menus, sliders, text fields, and panels that allow users to manipulate images, graphics, videos, and 3D models. For example, Photoshop provides consistent tools for adjustments like brightness, contrast, and hue, while Illustrator offers vector editing through pre-defined tools. Premiere Pro delivers video editing features through timelines and control panels, and Blender offers a complex interface for 3D modeling and animation. These tools are powerful yet rigid, requiring users to learn and adapt to the software’s structure. As a result, their complexity can overwhelm beginner users, and their interface, though it has undergone interactive user testing, may not always align with individual workflows or preferences, limiting the flexibility and personalization of the user experience.
2.2. Natural Language UIs
Using natural language as input to edit digital content has gained increasing interest due to its simplicity. This approach allows users to perform tasks like the generation and editing of images (OpenAI, 2024a; Runaway, 2024; Adobe, 2024a, d; Brooks et al., 2023), videos (Descript, 2024; Adobe, 2024f; Qin et al., 2024), data visualizations (Maddigan and Susnjak, 2023; Sah et al., 2024; Wang et al., 2023, 2024; Dibia, 2023), 3D models (Chang et al., 2017; Poole et al., 2022), and human motions (Tevet et al., 2022; Raab et al., 2023; Chen et al., 2023; Liu and Sra, 2024b, a) by simply describing the outputs they want rather than navigating complex software interfaces. Natural language UIs are transforming traditional workflows by making advanced editing more accessible to non-expert users, reducing the technical execution gulf, and enabling users to focus on creative intent. However, these UIs require articulation and can hurt usability as half the population is insufficiently articulate in writing to use natural language UIs well (Nielsen, 2023). Natural language UIs also struggle with precise control over specific details. They cannot effectively support interactive editing, which limits their ability to handle complex, fine-grained adjustments and provide prompt visual feedback of intermediate results typically required in iterative workflows.
2.3. Dynamically Generated UIs
@@
Recent research has studied generating dynamic UIs to provide users with precise, direct interactions for content editing. One prominent approach is the relaxation method, which involves creating UI widgets that provide generalized control over variables within a function or query. For instance, Mavo (Verou et al., 2016) allows users to construct interactive HTML pages by incorporating unique attributes, which, in turn, recommend editing widgets tailored to those attributes. Heer et al. (Heer et al., 2008) employ query relaxation techniques, allowing users to broaden their selection tasks effectively. Bespoke (Vaithilingam and Guo, 2019) utilizes user demonstrations to generate custom UIs for command-line applications. It applies rule-based heuristics to infer the semantic types of parameters in bash commands and thus dynamically create widgets for parameter manipulation. Another line of research, known as precision interfaces, uses SQL queries as a framework for UI generation (Zhang et al., 2018; Chen, 2020; Chen and Wu, 2022; Chen et al., 2022). These systems convert a sequence of input queries into interactive widgets, allowing users to adjust query parameters directly.
@@
In addition to generating UIs from command-line applications and SQL queries, recent advancements have focused on creating UIs from natural language. This approach allows simplicity by taking users’ natural language as input, and the system generates an interactive UI tailored to users’ task descriptions using LLM’s reasoning and code generation capabilities. For instance, DynaVis (Vaithilingam et al., 2024) generates dynamic interfaces for visualization editing based on natural language commands. By enabling users to describe changes to their visualizations through simple text, such as “rotate the x-labels by 30 degrees counter-clockwise”, DynaVis dynamically creates suitable widgets, such as a slider, for adjusting the rotation of x-labels. Similarly, Biscuit (Cheng et al., 2024) generates UIs to help users follow machine learning tutorials in Jupyter Notebooks using natural language queries. Users can describe tasks like “sample training data” or “adjust model hyperparameters”, and Biscuit generates UIs for the specified parameter adjustments. Although natural language-based UI generation supports both precision with interactable UI widgets and simplicity by taking natural language input, most prior work using standalone LLMs does not fully consider individual user preferences and domain-specific task requirements. This can result in generated interfaces that, while functional, may not be the most effective for specific users, especially when personalization and task-specific requirements are needed.
@@
Figure 1. Overview of the CrowdGenUI framework. This framework consists of four main components: it takes (1) user context as input, including task description and UI control preferences, and this information is passed to the LLM together with (2) a user preference library and (3) a collection of UI widget candidates. All the input data is used to guide the LLM reasoning process to obtain (4) generated UI widgets that are aligned with user preferences and task requirements.
@@
This figure is a flow chart that takes user context, user preference library, and UI widget candidates as input for LLMs, and generates UI as the output.
3. CrowdGenUI: A Framework that Aligns UI Generation with User Preference
3.1. Overview
Figure 1 provides an overview of our proposed framework, CrowdGenUI, which aligns generated UI widgets with task requirements and specific user preferences by incorporating a user preference library that captures users’ preferred UI widgets and their rationale to guide LLM-based UI generation.
Enhanced alignment through user preference integration: Unlike prior work on LLM-based UI generation, this framework integrates user preference insights directly into the LLM reasoning process. The user insights consist of crowdsourced preference data for UI widgets from real users. By designing a systematic reasoning technique to guide LLM reasoning, this framework is designed to boost LLM alignment in the UI generation application.
Flexibility and extendability through modular design: By using a modular-based architecture, this framework allows independent and flexible module updates to accommodate extendable contexts. This applies to all three input modules: (1) user context can evolve with updated task requirements and various preferences, ensuring continual alignment with specific user needs; (2) user preference library can continuously take updated user data, expanding the diversity and richness of user insights used during UI generation; (3) UI widget candidates can flexibly incorporate either static widget sets from established libraries or dynamically produced widgets, thereby expanding potential UI outcomes.
3.2. Usage Scenario
Before diving into the technical implementation details, we guide readers through an example usage scenario, shown in Figure 2, to demonstrate how CrowdGenUI can be used for a content editing task. Imagine that Alice wants to change the tone of an image from green, spring colors to warm, fall colors. She uses CrowdGenUI to edit her image through the following process:
Figure 2. Example usage scenario of CrowdGenUI. A user wants to change the color of an image to fall colors. They provide the image, describe the task, and specify their UI control preference. Taking user input, CrowdGenUI reasons the proper UI widgets and generates the code for widget implementation. The user uses the generated widgets to edit the image and can further update their input to obtain updated UIs, e.g., from efficient experiments using preset options to fine-grained control using a slider.
[…]
This figure demonstrates an example usage scenario using a flow chart. From left to right, the chart consists of interactive user input, UI widget reasoning and implementation, and content editing.
Step 1—Describe task and specify preferences: She describes her needs as “Change this image to fall colors” and provides the image she wants to edit. She also specifies that she prefers to experiment with image colors quickly with color tone overviews.
Step 2—Use CrowdGenUI for UI widget reasoning and implementation: CrowdGenUI processes Alice’s input and determines the most suitable UI widgets for her task. Since the task involves adjusting the image’s color tone efficiently, CrowdGenUI consults the crowdsourced user preference library to identify relevant tasks and commonly preferred UI widgets for efficiency. Based on this information, it reasons that preset buttons displaying color tone previews are the optimal choice. Next, it generates code to implement an image tone adjustment function and a preset button function, links the preset buttons to image tone adjustment, and returns the code for Alice to execute and obtain the widgets.
Step 3—Edit content with generated UI: Alice interacts with the preset buttons to explore various pre-defined tones. The color previews on buttons help her to anticipate how each selection will affect the image, and the preset values enable quick experimentation with simple button clicks.
Step 4—Update requirements for iterative UI generation: Although the preset buttons satisfy Alice’s initial needs, she finds an orange tone she likes but wants to make it slightly redder. Since the preset options do not allow that, she updates her request by asking for UI widgets that support finer tone adjustments. This brings her back to Step 2, where CrowdGenUI generates new widgets based on her updated requirements, and moves on to Step 3, where she continues refining her image using the newly generated widgets.
4. CrowdGenUI Implementation
Figure 3 illustrates a system prototype we implemented for our proposed framework. In this section, we dive into the technical details and implementation rationale.
Figure 3. System implementation pipeline. The system starts with natural language input specifying the task description and user preferences. With this input, the LLM applies chain-of-thought (CoT) reasoning to obtain reasoned widgets. Based on this reasoning process, the crowdsourced user preference library, containing the crowdsourcing tasks, user-preferred widgets, and their rationale, is integrated to augment widget reasoning. Additionally, the Jupyter Widget Library widget candidates are fed to the LLM for reasoned widget implementation. Lastly, the reasoned widgets and reference code snippets are passed to the LLM for code generation in Python. This allows widget implementation and interaction in Jupyter Notebook.
[…]
This figure is a flow chart showing the system implementation pipeline. Starting from the left, the task description, preference aspect, crowdsourced preference library, and UI widget candidates are passed as input to the LLM for CoT reasoning. As the reasoning output, reasoned widgets containing widget types and the corresponding rationale are returned. This output is further passed to the LLM for code generation with sample code as guidance. Finally, Python code is returned for widget display and interaction.
4.1. User Context
The system input is user context, which consists of task description and UI control preferences.
4.1.1. Task Description
The task description needs to specify the content to edit (e.g., an image), values to adjust (e.g., image color), and goals to achieve (e.g., changing the image to fall colors).
4.1.2. UI Control Preferences
UI control preferences can involve various preference aspects. In the current system, we use predictability, efficiency, and explorability to implement our prototype.
These preferences are chosen based on well-established UI design principles developed by prior research. Specifically, McKay’s UI is Communication (McKay, 2013) emphasizes that an intuitive interface relies on effective communication that balances multiple principles, including predictability, efficiency, discoverability, and explorability. McKay argues that intuitive UIs align with a user’s prior experiences, providing clear affordances and responsive feedback to guide interaction. Lidwell et al.’s Universal Principles of Design (Lidwell et al., 2010) expands on these ideas by highlighting fundamental design principles such as the flexibility-usability tradeoff, mapping, and mental models, which underscore the importance of aligning system controls with user expectations and simplifying complex interactions. Together, these principles highlight that successful UIs are predictable and efficient and allow for exploratory interaction without overwhelming the user, leading to our selection of predictability, efficiency, and explorability to guide LLM-based UI generation. We introduce the user preference library that we crowdsourced according to the three aspects in Section 5.
4.1.3. Input to the LLM
Apart from feeding the image to the LLM, the rest of the user context is represented as text input denoted by task-description ($T$) and preference-aspect ($A$) in JSON format.